# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ 1.6 Classification Safety and Training Stability
Safe training in classification tasks has been pursued through overconfidence mitigation, robustness to noise and perturbations, smooth decision boundaries, and algorithmic stability. Bousquet et al. [64] established that uniform stability provides generalization bounds, offering a theoretical foundation for safe objective functions and algorithm design. Hardt et al. [65] analyzed the data-dependent stability of SGD and formalized its relationship with learning rate, iteration count, and generalization. Guo et al. [66] highlighted the severe overconfidence problem in modern neural networks and demonstrated that temperature scaling can effectively calibrate predictions. Müller et al. [31] systematically showed that label smoothing improves both generalization and probabilistic calibration.
@@
In label noise settings, loss redesign has been a key approach. Patrini et al. [67] proposed correcting the loss by estimating the noise transition matrix, while Ghosh et al. [68] theoretically analyzed the robustness of symmetric losses. Zhang et al. [69] introduced a generalized loss that interpolates between cross-entropy and MAE, achieving a balance between robustness and convergence. Wang et al. [70] added a symmetric term to standard cross-entropy to enhance stability under mislabeled data.
@@
Methods that promote local smoothness and large margins help secure safe decision boundaries. Miyato et al. [71] enforced output invariance to small perturbations, ensuring local smoothness in classifiers. Zhang et al. [72] proposed mixup, which linearly interpolates both inputs and labels to smooth decision boundaries and reduce both overfitting and overconfidence. Thulasidasan et al. [73] showed that mixup is particularly effective for confidence calibration. Pascanu et al. [74] formalized that gradient clipping suppresses exploding updates and improves numerical stability during training. Loshchilov et al. [62] proposed AdamW to decouple weight decay from optimization steps, improving tuning and convergence stability.
@@
Safety against adversarial perturbations has been approached through robust optimization and certification methods. Madry et al. [75] established a standard min-max training procedure for learning classifiers robust to $\ell_{p}$ perturbations. Zhang et al. [76] proposed TRADES, characterizing the fundamental trade-off between robustness and accuracy and providing generalization bounds. Cissé et al. [77] proposed controlling Lipschitz bounds at the layer level to suppress perturbation amplification, and Cohen et al. [78] provided certified robustness within an $\ell_{2}$ radius via randomized smoothing.
@@ 1.7 NTK Extensions and Directional Regularization from Spectral and Kernel Perspectives
Jacot et al. [79] showed that as neural network width tends to infinity, the network linearizes into fixed kernel regression, with its learning dynamics governed by the Neural Tangent Kernel (NTK). Lee et al. [80] reinforced the generality of the NTK regime by proving that wide networks of arbitrary depth evolve like linear models under gradient descent. Chizat et al. [81] formalized the lazy training regime, where parameter changes remain small in the large-width limit, explaining kernel-like behavior when feature learning is limited. Du et al. [82] supported the optimization perspective of the NTK regime by proving that gradient descent converges to global minima under overparameterization.
The generalization of kernel regression has been shown to depend on the kernel eigenspectrum and task–model alignment. Bordelon et al. [83] quantified how learning curves vary with the kernel eigenspectrum and the target’s projection onto the kernel eigenfunctions. Canatar et al. [84] demonstrated that tasks aligned with the principal components of the kernel lead to higher sample efficiency, formalizing the alignment principle. Liang et al. [85] identified conditions under which interpolation without ridge regularization can still generalize, and Hastie et al. [22] analyzed the risk of high-dimensional minimum-norm interpolators, revealing the spectral dependence of kernel and linear solutions. Mei et al. [86] provided an exact high-dimensional characterization of the risk in random features regression, analyzing the spectral effects of kernel approximations.
From the kernel approximation viewpoint, Rahimi et al. [87] proposed random features as an efficient method for approximating kernels, and Rudi et al. [88] established generalization guarantees for such methods. Rahaman et al. [89] observed spectral bias in neural networks, showing that networks tend to learn functions from low to high frequencies, thereby linking neural network behavior to kernel and NTK-based analysis.
The classical framework for directional regularization has been established in Tikhonov regularization with weights and generalized ridge methods. Hoerl et al. [90] proposed shrinking along covariance directions to control estimation variance. Engl et al. [91] unified various regularization strategies—including Tikhonov and early stopping—under the spectral filtering perspective. Hansen [92] systematized how filter functions and penalty matrices can be used to selectively suppress or emphasize specific eigendirections.
@@ 2 Theoretical Analysis
This section aims to formalize AR as a reward term added to the empirical risk and to present the spectral safety condition and trust-region–based guarantees, the existence and convergence of solutions, conditions under which test error improves by correcting underfitting in small-sample regimes, the scaling rule for $\lambda$ as a function of sample size $n$, classification-specific issues such as the lack of a lower bound and remedies via trust-region or bounded margin rewards, the extension to the NTK regime of shallow MLPs (1–3 layers), and a stability-oriented interpretation. Detailed proofs are provided in the appendix.
2.1 Notation and Symbols
To ensure clarity, the following set of notations is fixed throughout the paper. Key symbols are defined precisely and used consistently in subsequent theorems and proofs without redundancy. Eigenvalues are denoted by $\sigma$, and the regularization coefficient is denoted by $\lambda$ to prevent notation conflicts.
@@ 2.1 Notation and Symbols
    • Training set: $S=\{(x_{i},y_{i})\}_{i=1}^{|S|}$
    • Parameters: $\theta$, predictor: $f_{\theta}$
@@
    • Regression: $\ell(\theta;x,y)=\tfrac{1}{2}(y-f_{\theta}(x))^{2}$
    • Classification: $\ell(\theta;x,y)=H(p_{y},q_{\theta}(x))$
@@
  • Linear Regression Setup
    • Design matrix: $X\in\mathbb{R}^{|S|\times p}$
    • Sample covariance: $\hat{\Sigma}=\tfrac{1}{|S|}X^{\top}X$
    • Eigenvalue notation: $\sigma_{\min}(\cdot),\ \sigma_{\max}(\cdot)$
@@
    • Smoother: $S_{\lambda}=X(X^{\top}X-|S|\,\lambda I)^{-1}X^{\top}$
    • Degrees of freedom: $\operatorname{tr}(S_{\lambda})$
@@
  • Classification-Specific Symbols
    • Margin: $m_{\theta}(x,y)$
    • Bounded margin reward function: $\phi$
    • Network Lipschitz constant: $G$
    • Margin threshold: $\gamma$
@@
    • Kernel: $K$
@@ 2.2 Safety (Existence and Convergence)
Theorem 1 (Spectral Safety Condition and Strong Convexity in Regression).
@@
The Hessian is given by $\nabla^{2}\hat{F}_{\lambda}=\hat{\Sigma}-\lambda I$. The smallest eigenvalue is $\sigma_{\min}(\hat{\Sigma})-\lambda>0$, which ensures strong convexity. ∎
Corollary 2 (Closed-form Solution and Gradient Descent Stability).
The unique minimizer is given by $\hat{\theta}_{\lambda}=(X^{\top}X-|S|\,\lambda I)^{-1}X^{\top}y$. Moreover, if $0<\eta<\tfrac{2}{\sigma_{\max}(\hat{\Sigma})-\lambda}$, then gradient descent converges linearly.
Lemma 3 (Lower Bound Guarantee in Classification).
In classification, the objective $H-\lambda\|\theta\|^{2}$ may be unbounded below on separable data. A minimizer exists if one of the following conditions holds. First, optimization is performed under a trust-region constraint $\|\theta\|\leq B$. Second, a bounded margin reward function $\phi$ is used, yielding:
$\hat{F}_{\lambda}(\theta)=\frac{1}{|S|}\sum H\big{(}p_{y},q_{\theta}(x)\big{)}-\lambda\cdot\frac{1}{|S|}\sum\phi\big{(}m_{\theta}(x,y)\big{)}$
@@
A trust-region constraint defines a compact domain, which ensures that the continuous objective admits a minimum. The bounded reward $\phi$ provides a lower bound for the objective. ∎
@@ 2.3 Degrees of Freedom and Sample-size–Dependent Scaling (Regression)
The degrees of freedom of the linear smoother $S_{\lambda}$ is given by $\operatorname{tr}(S_{\lambda})=\sum_{j=1}^{p}\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}$, where $\{\sigma_{j}\}$ are the eigenvalues of $\hat{\Sigma}$. This quantity increases monotonically in $\lambda$ within the spectral safety condition, and amplifies directions with smaller $\sigma_{j}$ more strongly. Under the generative model $y=X\theta^{\star}+\varepsilon$, the following optimism identity holds:
@@
Subtracting the two optimism identities yields the inequality as a sufficient condition for test risk improvement. ∎
2.4 Sample-size Calibration Rule (Regression)
DoF targeting. For $\kappa\gtrsim 1$, solve $\operatorname{tr}(S_{\lambda})/|S|=\kappa$ using monotonicity to uniquely determine $\lambda$ such that per-sample complexity remains approximately constant.
Power-law decay schedule.
$|\lambda(n)|=|\lambda(n_{0})|\left(\frac{n_{0}}{n}\right)^{\alpha},\qquad\alpha=1\ \text{or more conservatively }\alpha>1.$
2.5 Spectral Weighting and Control over Noise Directions
Let $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, with $\hat{\Sigma}=U\,\mathrm{diag}(\sigma_{j})\,U^{\top}$ and $W=U\,\mathrm{diag}(w_{j})\,U^{\top}$. Then,
@@ 2.5 Spectral Weighting and Control over Noise Directions
which allows mitigation of over-amplification in low-$\sigma_{j}$ directions by assigning smaller $w_{j}$.
2.6 Classification: Bounded Margin Reward and Generalization
If the network has Lipschitz constant $G$ and the reward function $\phi$ is both bounded and Lipschitz, then the standard margin-based generalization bound implies:
@@ 2.6 Classification: Bounded Margin Reward and Generalization
Theorem 5 (Sample-size Scaling in Classification).
Let $|\lambda(n)|=|\lambda(n_{0})|\left(\tfrac{n_{0}}{n}\right)^{\alpha}$ for $\alpha=\tfrac{1}{2}$ or more conservatively $\alpha>\tfrac{1}{2}$. Then, as $n$ increases, the schedule ensures that $|\lambda(n)|\to 0$ while maintaining balance between empirical risk and complexity terms.
@@
Since the dominant order of the complexity term is $O(1/\sqrt{n})$, setting $|\lambda|$ to decay at the same rate maintains the bound’s balance. ∎

### v2

@@ 2.4 Safety in Classification and Training Stability
Safe learning in classification has been pursued through the suppression of overconfidence, robustness to noise and perturbations, smooth decision boundaries, and algorithmic stability. Bousquet et al. [64] showed that uniform stability provides generalization bounds, establishing theoretical criteria for safe objective and algorithm design, while Hardt et al. [65] analyzed the data-dependent stability of SGD, formalizing its connection to learning rates, iteration counts, and generalization. Guo et al. [66] identified the severe overconfidence problem in modern neural networks and demonstrated that temperature scaling can effectively calibrate predictions. Müller et al. [31] systematically reviewed label smoothing, highlighting its benefits for both generalization and probability calibration.
@@
In noisy-label settings, redesigning the loss function has been a primary approach. Patrini et al. [67] proposed estimating the noise transition matrix to correct losses, and Ghosh et al. [68] theoretically showed that symmetric losses are robust against label noise. Zhang et al. [69] introduced a generalized loss interpolating between cross-entropy and MAE to combine robustness with convergence, while Wang et al. [70] improved stability on mislabeled data by augmenting standard cross-entropy with a symmetric component.
@@
Methods that encourage local smoothness and large margins enhance the safety of decision boundaries. Miyato et al. [71] optimized for output invariance against small perturbations to guarantee local smoothness of classifiers, and Zhang et al. [72] promoted smooth boundaries via linear interpolation of inputs and labels, thereby reducing both overfitting and overconfidence. Thulasidasan et al. [73] showed that mixup is particularly effective in improving classification calibration. Pascanu et al. [74] formalized that gradient clipping suppresses exploding gradients, enhancing numerical stability during training, while Loshchilov et al. [62] improved tuning and convergence stability by decoupling weight decay from the optimization step in AdamW.
@@
Safety against adversarial perturbations has been addressed through robust optimization and certification methods. Madry et al. [75] established minimax training as a standard procedure for learning classifiers robust to $\ell_{p}$ perturbations. Zhang et al. [76] proposed TRADES, which clarified the principled trade-off between robustness and accuracy and provided generalization bounds. Cissé et al. [77] proposed directly controlling Lipschitz bounds of layers to suppress perturbation amplification, while Cohen et al. [78] introduced randomized smoothing to provide certified robustness within an $\ell_{2}$ radius.
@@ 2.5 NTK Extensions and Directional Regularization from Spectral and Kernel Perspectives
Jacot et al. [79] showed that as network width tends to infinity, neural networks linearize into fixed kernel regression, with learning dynamics governed by the Neural Tangent Kernel (NTK). Lee et al. [80] strengthened the universality of the NTK regime by proving that wide networks of arbitrary depth evolve like linear models under gradient descent. Chizat et al. [81] formalized the lazy training regime, in which parameter updates become vanishingly small at large widths, explaining kernel-like behavior when feature learning is limited. Du et al. [82] provided optimization support for the NTK regime by showing that gradient descent finds global minima under overparameterization.
Generalization in kernel regression has been analyzed as depending on the eigenspectrum of the kernel and its alignment with the target function. Bordelon et al. [83] quantified how learning curves vary with kernel eigen-spectra and the target’s eigenfunction components. Canatar et al. [84] established the alignment principle, showing that sample efficiency improves when tasks align with principal kernel directions. Liang et al. [85] provided conditions under which kernel interpolation regression generalizes without ridge regularization, while Hastie et al. [22] analyzed the risk of high-dimensional minimum-norm interpolators, revealing spectral dependencies of both kernel and linear solutions. Mei et al. [86] exactly characterized the risk of random feature regression in high dimensions, thereby analyzing the spectral effects of kernel approximations.
From the perspective of kernel approximation, Rahimi et al. [87] proposed random features as an efficient procedure for approximating kernels, and Rudi et al. [88] systematized the generalization guarantees of such approximations. Rahaman et al. [89] observed the spectral bias of neural networks, whereby functions are learned progressively from low to high frequencies, providing a link to kernel- and NTK-based analyses.
The classical framework of directional regularization has been established in weighted Tikhonov regularization and generalized ridge methods. Hoerl et al. [90] introduced criteria for controlling estimation variance through shrinkage along covariance directions. Engl et al. [91] organized equivalences among Tikhonov, early stopping, and other regularization methods from the perspective of spectral filtering. Hansen [92] systematized procedures for selectively suppressing or amplifying specific eigen-directions through filter functions and penalty matrices.
@@ 3 Theoretical analysis
AR을 경험위험에 대한 보상 항으로 정식화하고, 안전구간과 해의 존재 및 수렴, 작은 표본에서 언더핏을 교정할 때의 테스트 오류 개선 조건, 표본 수 nn에 따른 $\lambda$의 스케일 규칙, 분류에서의 하한 부재 문제와 이를 해결하는 신뢰영역 또는 상계 마진 보상, 얕은 MLP 1–3층의 NTK 레짐 확장, 그리고 안정성 관점을 제시한다. 상세 증명은 부록에 수록한다.
3.1 Notation
For clarity of exposition, we fix the following set of symbols to be used consistently throughout the paper. Core notations are explicitly defined here and will be used without redundancy in subsequent theorems and proofs. In particular, eigenvalues are denoted by $\sigma$, and regularization coefficients by $\lambda$, to avoid conflicts.
@@ 3.1 Notation
    • Training set $S=\{(x_{i},y_{i})\}_{i=1}^{|S|}$
    • Parameters $\theta$, predictor $f_{\theta}$
@@
    • Regression: $\ell(\theta;x,y)=\tfrac{1}{2}\big{(}y-f_{\theta}(x)\big{)}^{2}$
    • Classification: $\ell(\theta;x,y)=H\!\big{(}p_{y},\,q_{\theta}(x)\big{)}$
@@
  • Linear regression setting
    • Design matrix $X\in\mathbb{R}^{|S|\times p}$
    • Sample covariance $\hat{\Sigma}=\tfrac{1}{|S|}X^{\top}X$
    • Eigenvalue notation $\sigma_{\min}(\cdot),\ \sigma_{\max}(\cdot)$
@@
    • Smoother $S_{\lambda}=X\big{(}X^{\top}X-|S|\,\lambda I\big{)}^{-1}X^{\top}$
    • Degrees of freedom $\operatorname{tr}(S_{\lambda})$
@@
  • Classification-related notation
    • Margin $m_{\theta}(x,y)$
    • Bounded margin reward $\phi$
    • Network Lipschitz constant $G$
    • Margin threshold $\gamma$
@@
    • Kernel $K$
Assumption (applicability in linear regression).
[…]
@@ 3.2 Safety (Existence and Convergence)
Theorem 1 (Safe region and strong convexity in regression).
@@
The Hessian is $\nabla^{2}\hat{F}_{\lambda}=\hat{\Sigma}-\lambda I$. If its minimum eigenvalue satisfies $\sigma_{\min}(\hat{\Sigma})-\lambda>0$, strong convexity holds. ∎
Remark 1 (Applicability).
[…]
Corollary 2 (Closed-form solution and stability of gradient descent).
The unique solution is given by $\hat{\theta}_{\lambda}=(X^{\top}X-|S|\,\lambda I)^{-1}X^{\top}y$. Moreover, if $0<\eta<\tfrac{2}{\sigma_{\max}(\hat{\Sigma})-\lambda}$, gradient descent converges linearly.
Lemma 3 (Classification: boundedness from below and potential non-attainment).
Consider separable data and an upper-bounded margin reward $\phi:\mathbb{R}\to[0,\phi_{\max}]$:
$\hat{F}^{(\phi)}_{\lambda}(\theta)=\frac{1}{|S|}\sum H\big{(}p_{y},q_{\theta}(x)\big{)}\;-\;\lambda\cdot\frac{1}{|S|}\sum\phi\!\big{(}m_{\theta}(x,y)\big{)},\quad\lambda>0.$
For all $\theta$, we have
$\hat{F}^{(\phi)}_{\lambda}(\theta)\;\geq\;-\,\lambda\,\phi_{\max},$
so the objective is bounded below. However, in non-compact domains with separable data, $\inf_{\theta}\hat{F}^{(\phi)}_{\lambda}(\theta)=-\lambda\,\phi_{\max}$ may only be approached as $\|\theta\|\to\infty$, and the minimum may not be attained.
@@
Since $H\geq 0$ and $\phi\leq\phi_{\max}$, the lower bound follows immediately. For separable data, consider the trajectory $\theta_{t}=t\,u$. Then $H\to 0$ and $m_{\theta_{t}}\!\to\infty$, so $\phi(m_{\theta_{t}})\!\to\phi_{\max}$. However, since $\|\theta_{t}\|\!\to\infty$, only the infimum is achieved, and the minimum may not exist. ∎
Lemma 4 (Classification: conditions for existence of minimizers).
When optimizing over a trust region $\Theta_{B}=\{\theta:\|\theta\|\leq B\}$, the continuous objective $\hat{F}^{(\phi)}_{\lambda}$ admits a minimizer by the Weierstrass theorem. Even in the unconstrained case, adding a small positive regularization $\varepsilon>0$ and defining
$\hat{F}_{\lambda,\varepsilon}(\theta):=\hat{F}^{(\phi)}_{\lambda}(\theta)\;+\;\varepsilon\|\theta\|^{2}$
makes the objective coercive, guaranteeing the existence of a minimizer. Alternatively, parameter or logits can be suitably clipped so that all sublevel sets are compact, in which case a minimizer likewise exists.
Proof sketch.
Since $\Theta_{B}$ is compact, continuity of $\hat{F}^{(\phi)}_{\lambda}$ ensures the existence of a minimum by the Weierstrass theorem. With added regularization, $\hat{F}_{\lambda,\varepsilon}(\theta)\geq-\lambda\phi_{\max}+\varepsilon\|\theta\|^{2}\to+\infty$, so the objective is coercive, and lower semicontinuity guarantees a minimizer. Finally, clipping parameters or logits renders sublevel sets compact, and the same reasoning ensures existence of minimizers. ∎
@@ 3.3 Degrees of Freedom and Sample-size–dependent Scaling (Regression)
For the linear smoother $S_{\lambda}$, the degrees of freedom are given by
$\operatorname{tr}(S_{\lambda})=\sum_{j=1}^{p}\frac{\sigma_{j}}{\sigma_{j}-\lambda},$
where $\{\sigma_{j}\}$ are the eigenvalues of $\hat{\Sigma}$. Within the safe region, this quantity is monotone increasing in $\lambda$, amplifying directions associated with small $\sigma_{j}$. In the generative model $y=X\theta^{\star}+\varepsilon$, the following optimism identity holds:
@@
Taking the difference of the optimism identities yields the stated inequality as a sufficient condition for reduction of test risk. ∎
3.4 Sample-size Calibration Rules (Regression)
Degrees-of-freedom targeting: For $\kappa\gtrsim 1$, one uniquely solves $\operatorname{tr}(S_{\lambda})/|S|=\kappa$ by monotonicity, thereby maintaining constant per-sample complexity.
Power-law schedule:
$|\lambda(n)|=|\lambda(n_{0})|\Big{(}\frac{n_{0}}{n}\Big{)}^{\alpha},\qquad\alpha=1\ \text{or, more conservatively, }\alpha>1.$
3.5 Spectral Weighting and Control of Noise Directions
If $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, $\hat{\Sigma}=U\operatorname{diag}(\sigma_{j})U^{\top}$, and $W=U\operatorname{diag}(w_{j})U^{\top}$, then
@@ 3.5 Spectral Weighting and Control of Noise Directions
so that assigning small $w_{j}$ to small $\sigma_{j}$ can suppress excessive amplification.
Remark 2 (Basis alignment for spectral weighting).
The closed-form expansion $\operatorname{tr}\!\big{(}S^{(W)}_{\lambda}\big{)}=\sum_{j}\tfrac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}}$ is derived when $W$ and $\widehat{\Sigma}$ are simultaneously diagonalizable in the same eigenbasis ($\widehat{\Sigma}=U\operatorname{diag}(\sigma_{j})U^{\top},\;W=U\operatorname{diag}(w_{j})U^{\top}$). For general $W$, it always holds that
$\operatorname{tr}\!\big{(}S^{(W)}_{\lambda}\big{)}=\operatorname{tr}\!\big{(}\widehat{\Sigma}(\widehat{\Sigma}-\lambda W)^{-1}\big{)},$
but the eigenvalue-sum form may not directly apply. In practice, it is safer to design $W$ aligned with the eigenbasis of $\widehat{\Sigma}$, or alternatively, to define $A=\widehat{\Sigma}^{-1/2}W\widehat{\Sigma}^{-1/2}$ via whitening and impose the safety condition $\lambda<1/\sigma_{\max}(A)$ to ensure stability.
3.6 Classification: Bounded Margin Rewards and Generalization
Suppose the network has Lipschitz constant $G$ and the reward function $\phi$ is both bounded and Lipschitz. Then, from standard margin-based bounds, we obtain
@@ 3.6 Classification: Bounded Margin Rewards and Generalization
Theorem 6 (Sample-dependent scheduling in classification).
In the classification setting, assume the loss $H$ is $L_{\ell}$-Lipschitz with respect to the logits, and the margin reward $\phi$ satisfies $0\leq\phi\leq\phi_{\max}$ and is $L_{\phi}$-Lipschitz. Further, suppose the sensitivity of the margin to parameter changes is bounded as $\|\nabla_{\theta}m_{\theta}(x,y)\|\leq L_{m}$ due to trust-region or clipping constraints, and that the effective strong convexity satisfies $\mu_{\mathrm{eff}}(n)=\mu-|\,\lambda(n)\,|\,\alpha_{R}\geq\delta\,\mu>0$, which requires $|\,\lambda(n)\,|\leq(1-\delta)\mu/\alpha_{R}$. Then there exist constants $C_{1},C_{2}>0$ such that
$\mathbb{E}\big{[}\mathcal{R}(\hat{\theta}_{\lambda(n)})\big{]}-\mathcal{R}^{*}\;\;\leq\;\;C_{1}\,|\,\lambda(n)\,|\;+\;\frac{C_{2}}{\sqrt{n}}.$
Thus, if we set $|\,\lambda(n)\,|=|\,\lambda(n_{0})\,|\big{(}\tfrac{n_{0}}{n}\big{)}^{\alpha}$ with $\alpha\geq\tfrac{1}{2}$, then
$\mathbb{E}\big{[}\mathcal{R}(\hat{\theta}_{\lambda(n)})\big{]}-\mathcal{R}^{*}\;\;\leq\;\;\frac{C_{1}\,|\,\lambda(n_{0})\,|\,n_{0}^{\alpha}+C_{2}}{\sqrt{n}}.$
In particular, for $\alpha=\tfrac{1}{2}$, the two terms are balanced at the same rate, whereas for $\alpha>\tfrac{1}{2}$, the reward bias decays faster and the complexity term dominates. Moreover, as $n\to\infty$, we have $|\,\lambda(n)\,|\to 0$.
@@
If the effective strong convexity satisfies $\mu_{\mathrm{eff}}(n)\geq\delta\mu>0$, then uniform stability (or margin- and Rademacher-based complexity bounds) yield a generalization gap of order $O(1/\sqrt{n})$. On the other hand, the reward component of the objective is given by the sample average $R(\theta)=\tfrac{1}{n}\sum\phi(m_{\theta})$, whose parameter Lipschitz constant is bounded by $L_{\phi}L_{m}$. Thus, an additional bias term proportional to the reward strength, $C_{1}|\,\lambda(n)\,|$, appears. Combining both contributions yields the stated bound. The final inequality follows by substituting $|\,\lambda(n)\,|=|\,\lambda(n_{0})\,|(n_{0}/n)^{\alpha}$. ∎